At an air show a jet flies directly toward the stands at a speed of 1200km/h, emitting a frequency of 3500Hz, on a day when the speed of sound is 342m/s. What is the frequency received by stationary observers?
Answer #1

& = 1200 km/u = 333•33 m/s sound = 342 m/s f - 10. Vsound - source ource = 350ool 342 342 - 333.33 ) = 1381153.8461 |- |- 3 8